# Reliability and Interpretability in Science and Deep Learning

### v2

@@ Abstract
In recent years, the question of the reliability of Machine Learning (ML) methods has acquired significant importance, and the analysis of the associated uncertainties has motivated a growing amount of research. However, most of these studies have applied standard error analysis to ML models—and in particular Deep Neural Network (DNN) models—which represent a rather significant departure from standard scientific modelling. It is therefore necessary to integrate the standard error analysis with a deeper epistemological analysis of the possible differences between DNN models and standard scientific modelling and the possible implications of these differences in the assessment of reliability. This article offers several contributions. First, it emphasises the ubiquitous role of model assumptions (both in ML and traditional Science) against the illusion of theory-free science. Secondly, model assumptions are analysed from the point of view of their (epistemic) complexity, which is shown to be language-independent. It is argued that the high epistemic complexity of DNN models hinders the estimate of their reliability and also their prospect of long term progress. Some potential ways forward are suggested. Thirdly, this article identifies the close relation between a model’s epistemic complexity and its interpretability, as introduced in the context of responsible AI. This clarifies in which sense—and to what extent—the lack of understanding of a model (black-box problem) impacts its interpretability in a way that is independent of individual skills. It also clarifies how interpretability is a precondition for assessing the reliability of any model, which cannot be based on statistical analysis alone. This article focuses on the comparison between traditional scientific models and DNN models. However, Random Forest (RF) and Logistic Regression (LR) models are also briefly considered.
@@ 1 Introduction
First, we cannot completely exclude the possibility that successful DNN algorithms are over-fitting the collections of datasets that are used to train and test them (Recht et al., 2018). In fact, because of the difficulty of collecting good quality labelled data, very few very popular datasets (in particular CIFAR-10, Krizhevsky et al. (2014), ImageNet, Russakovsky et al. (2015) and a few others) represent the unique benchmark for the majority of the research work on DNN. This violates one of the key statistical assumptions of all ML methods, namely that the parameters are set independently of the test data. Strictly speaking, if some data have been used in an article we read before designing our new DNN architecture, the same data should not be used to test our new architecture. In practice, it can be hard to adhere to this requirement.
Secondly, successful applications of ML methods are much more likely to be published than unsuccessful ones. While publication biases (Dickersin et al., 1987) affect all types of scientific research, one can expect a much bigger impact on ML research, whose persuasive power relies crucially on its empirical success. This is related to the lack of full understanding of the mechanism that allows DNN to learn some features and ignore others.
Furthermore, assessment of confidence levels of ML predictions is notoriously difficult, especially in the case of DNN. The best evidence of these difficulties is provided by so-called adversarial examples (Szegedy et al., 2014): images that are misclassified with very high confidence by the DNN classifier, although they are humanly indistinguishable from other images that are correctly classified. Adversarial examples are much more pervasive in DNN than initially expected (Carlini and Wagner, 2017). The problem is not the occurrence of misclassification itself, but the high sensitivity to invisible features and the very high confidence (easily $>99\%$) quoted by the classifier. This is clearly not a reliable error estimate [^1] It also does not help to say that adversarial examples have negligible probability among natural images, because such statement cannot be made quantitative, as we lack a suitable characterisation of the space of possible natural images. Moreover, adversarials have been also reproduced from natural images (Hendrycks et al., 2019)..
@@
ML methods in general, and DNN in particular, are being used successfully in many contexts where assessing their expected errors is not necessary. This applies, for example, to all those cases where the ML method improves the search efficiency of candidate solutions that can be subsequently verified by other means (Duede, 2023). This is the case in a vast range of applications spanning material discovery, drug discovery, predictive maintenance, fraud detection (Baesens et al., 2015), code suggestion (Brown et al., 2020), protein folding (Jumper et al., 2021) and many others. In those cases, assessing the reliability is the responsibility of the independent check, while the ML method merely improves the efficiency of the overall process. However, there are also applications where independent checks are not practical (e.g. for safety-critical real-time systems (Buttazzo, 2022)) or cases where it is crucial to estimate the amount of missed solutions (e.g. compliance applications). In those cases, assessing the reliability of the ML method is very important.
This paper addresses the reliability of DNN methods from a fundamental epistemological point of view. It is very important to combine this perspective with a purely statistical one because even traditional science does not offer an absolute (assumption-free) guarantee that its predictions have specific probabilities. So, we must understand to what extent DNN models rest on similar grounds as those employed by more traditional scientific disciplines and to what extent they might suffer from fundamentally different problems. It is also worth noting that the DNN approach displays many aspects of a new paradigm (Kuhn, 1962) with respect to traditional scientific disciplines. While this paper argues that the comparison with traditional science is possible and necessary, it certainly cannot be performed naively.
@@
The topic of this article falls at the intersection of the theory of complexity (Li and Vitányi, 2019), DNNs (Goodfellow et al., 2016; LeCun et al., 2015), responsible AI (MacIntyre et al., 2021; Eitel-Porter, 2021) and the epistemology of ML, which has attracted considerable attention recently (see e.g. Desai et al. (2022); Leonelli (2020)). As opposed to many works on this topic (but similar to Kasirzadeh et al. (2023); Zenil and Bringsjord (2020) among others), this article stresses the important peculiarities of DNNs within ML methods. The present approach has some similarities with the one adopted in (Watson and Floridi, 2021). Some of the most important differences are: (a) the present focus is on reliability, that naturally leads to consider global, rather than local interpretability; (b) for the same reason, the relevance defined in (Watson and Floridi, 2021) is less applicable and it is not considered here; (c) finally, Watson and Floridi (2021) consider a subjective notion of simplicity, while we focus on the hardcore complexity that no human can reduce, regardless of language and individual skills. This article investigates the objective foundations of reliability assessment (i.e. based on well-defined assumptions). Hence, it is largely unrelated to the literature that investigates the sociological basis for the trust in ML models, or analogies between ML and human behaviours (see e.g., Duede (2022); Clark and Khosrowi (2022); Tamir and Shech (2023)). Further comparisons with the literature are provided in the main text.
@@ 2 Assessing the reliability of model predictions in science and ML
To assess the reliability of any model we must be able to estimate the uncertainty of its predictions, in some precise and useful probabilistic sense. To start, it is convenient to distinguish between statistical and systematic uncertainties. Traditionally, statistical uncertainties are defined as those error sources which have a known statistical distribution (Bohm and Zech, 2017) [^2] The ML literature adopts sometimes similar but slightly different classification between statistical (aka random, aleatoric or ‘data’) uncertainties on one side and systematic (aka epistemic or model) uncertainties on the other (Hüllermeier and Waegeman, 2021; Gawlikowski et al., 2021). Here we stick to the traditional one, because it clarifies the role of the model assumptions, which is crucial in our discussion.. These uncertainties can be safely analysed via statistical methods. Systematic uncertainties are all the others: they may stem from systematic distortions in the measurement devices, from imbalances in the data selection, from inaccurate models, from approximation errors, or from inaccurate parameter fitting. As emphasised in Bohm and Zech (2017), even random noise with unknown statistical properties must be classified as a systematic effect [^3] Note that probability distributions without finite mean and variance, for which the central limit theorem does not apply, cannot be characterised empirically in most cases. Given the ubiquity of, e.g., self-similar phenomena (Barenblatt, 1996) it is clear that one cannot certainly assume that every noise generates statistical uncertainties.. Statistical uncertainties can be very difficult to estimate in practice, but the process is conceptually clear. They are knwon-unknown, as they are expected by the model. Systematic uncertainties are unknown-unknown and they can be estimated only by enriching the model assumptions. This raises the question of which assumptions are acceptable to assess our model assumptions?
@@ 2.1 Sources of errors
It is useful to distinguish four different sources of errors and see how they affect ML and TS models in different ways. Errors may stem (a) from data measurements, including the labels of training data; (b) from the model, which may not faithfully describe the actual phenomena in scope; (c) from any approximation that may be applied to the model to derive further predictions; and (d) from fitting the parameters that are left free by the model [^5] Type (a) and (d) may have both statistical and systematic contributions, while type (b) and (c) are purely systematic..
@@
Sources (b) and (d) are the most interesting ones when comparing TS and ML models. Model errors (b) can be reduced by extending the basic (ideal) model with further elements designed to model the deviations from the basic model. However, this enrichment comes at the expense of more assumptions and more parameters to be fitted (generating more uncertainties of type (d)). We may say that unknown model errors can be traded for more model assumptions and more parameters.
Source (d) refers to errors in the determination of the optimal parameters, even when the model might provide a correct representation of the phenomena. This may happen because of limited data, noisy data (source (a)), an inaccurate minimisation algorithm or an inexact parameter fitting process. Error source (d) has a statistical component, which is propagated from noisy data (a) and can be analysed statistically (along the lines of e.g. Marshall and Blencoe (2005) and references therein). However, source (d) also has systematic components, due to the complexity of determining the optimal parameters and also due to the unknown statistical properties of the data666for some classes of noise sources, collecting sufficient data can be completely hopeless in practice (see, e.g., de Forcrand (2009)) or even in principle (see e.g., Gabaix (2009)).. Moreover, increasing the number of parameters requires, in general, an exponential increase in the amount of data required for their determination [^7] An exponential increase is expected, in the generic case, according to the curse of dimensionality (Giraud, 2021).. In conclusion, also estimating (d) requires making suitable assumptions.
@@ 2.2 The illusion of predictions without assumptions
The simple reason why this idea is fallacious is the well-known (but often overlooked) underdetermination of the theory by the data (Duhem, 1954; Quine, 1975; Stanford, 2021). No matter how much data we have collected, there are always infinite different models that fit those data within any given approximation. Moreover, for every conceivable prediction, there exist infinite models that still fit all the previous data and also deliver the desired prediction [^8] In the context of the classical analysis of knowledge, the same problem is usually referred as the Gettier problem (see e.g. Floridi (2004) and references therein).. Data alone cannot enforce any conclusion.
@@
Another frequent fallacy is the idea that we should only accept assumptions that have been tested. But how much and under which conditions should they be tested? Newton’s laws were tested extensively under all conceivable circumstances for over two centuries before discovering that they do not always hold. There is no way to fully test our assumptions (or to define their scope of applicability) in a future-proof way. They may fail in contexts that we cannot formulate because we cannot foresee them.
@@ 3.1 Bayesian Error Estimates
However, such an argument is based on the unrealistic idea that the model is fixed, while we can rather easily collect more data. This is not what usually happens in real applications. In most cases, instead, collecting valuable data can be very hard, while producing new models is much easier. This is the case in traditional science, where planning a new experiment might take decades, while new models are published every day. It is the case also in most ML applications, where collecting labelled data can be extremely labour intensive, while new ML models are produced at the rate of millions per second by the ML algorithm. Under these conditions, there is no guarantee that the dependency on the priors will be sufficiently suppressed by the time the model is used.
@@
Increasing the number of parameters might look like a way to eliminate the need of changing the model. But, the Doob and Berstein - von Mises theorems do not apply when the number of parameters is of the same order as the size of the dataset or more (Johnstone, 2010; Sims, 2010). Extensions exist for high-dimensional problems (Banerjee et al., 2021) and even infinite-dimensional (aka nonparametric) ones (Ghosal and Van der Vaart, 2017). However, those extensions rely on the assumption that most parameters are very close to zero, which is achieved via LASSO priors or equivalent suppression mechanisms (Banerjee et al., 2021). Note that this is a much stronger assumption than requiring that all the relevant parameters live in some low-dimensional subspace of a high-dimensional model. A parameter suppression mechanism imposes a sparse realisation of the given parametrisation. As a result, even in the asymptotic limit of infinite statistics, the posteriors depend on the choice of the parametrisation, similar to the case of low-dimensional models. In other words, nonparametric models are not a way to eliminate model dependency. The Bayesian approach is an excellent framework to discuss the probability of a model within a given class of models. But the universe of all possible classes of models is really big and the Bayesian approach is not a way to estimate the probability of those overall model assumptions. This holds for both ML and TS models.
@@ 3.4 Predictions are NOT all you need
Can’t we just use the success rate of the test dataset as an estimate of the expected prediction error? Intuitively, one expects that novel predictions on unseen data do provide a meaningful test. However, formulating this idea precisely is fraught with riddles (Crupi, 2021). In fact, this intuitive expectation is based on some hidden assumptions. In particular, we are implicitly assuming that the distribution of the data will not change significantly. But this is not a useful assumption, because we cannot deduce anything from the fact that “the distribution of the data does not change” without assuming anything about the distribution itself.
This is obviously a problem for DNN models, but also for TS models. In fact, as already remarked, we cannot define the scope of applicability of any theory in a way that is future-proof: the model may fail in circumstances that we currently cannot even imagine. But for DNNs the problem is somewhat harder because the domain of applicability is difficult to define also with reference to features that we do know.
In the philosophy of science, the idea that “predictions are all you need” goes under the name of strong predictivism (Barnes, 2022). However, the strong version of predictivism has essentially no support and it is used only as a target for criticisms. In fact, predictions alone—no matter how impressive—are never sufficient to warrant trust in a model (Barnes, 2022; Votsis et al., 2014). Trustworthy predictions require trustworthy premises. Imagine that a crazy newly proposed theory predicts, correctly, something as unlikely as the exact outcome of a national lottery (and otherwise agrees with standard theories). Would we feel compelled to embrace the new theory? Certainly not. Likelihood by itself is never sufficient to determine theory selection. Other non-empirical aspects are also important. But which ones?
This problem applies also to TS models, but it becomes really disturbing for DNN models whose persuasive power relies almost entirely on their empirical success. Although DNN successes are impressive, it is impossible to derive any quantitative level of confirmation from them without additional assumptions. For example, published results certainly suffer a positive selection. Kaggle competitions (Goldbloom, 2015) are very interesting also from a philosophical point of view because they provide a controlled framework for assessment. However, they also cannot be used directly to confirm either general DNN architectures or specific DNN models, because submissions are strongly biased in terms of the approach used and the competencies of the participants. The fact that some successful DNN architectures keep being successful over time is certainly convincing evidence that they can learn something valuable. However, also in those cases, we cannot exclude that the DNNs have actually learned unwanted features that happened to have a strong spurious correlation with the labels (see, e.g., Xiao et al. (2021) for an example of this phenomenon). This confirms that also DNN predictions are only valuable when they stem from a valuable model. What characterises a valuable model, both in TS and in ML? This question is addressed in more depth in Sec. 4.

### v3

@@ Abstract
In recent years, the question of the reliability of Machine Learning (ML) methods has acquired significant importance, and the analysis of the associated uncertainties has motivated a growing amount of research. However, most of these studies have applied standard error analysis to ML models—and in particular Deep Neural Network (DNN) models—which represent a rather significant departure from standard scientific modelling. It is therefore necessary to integrate the standard error analysis with a deeper epistemological analysis of the possible differences between DNN models and standard scientific modelling and the possible implications of these differences in the assessment of reliability. This article offers several contributions. First, it emphasises the ubiquitous role of model assumptions (both in ML and traditional science) against the illusion of theory-free science. Secondly, model assumptions are analysed from the point of view of their (epistemic) complexity, which is shown to be language-independent. It is argued that the high epistemic complexity of DNN models hinders the estimate of their reliability and also their prospect of long term progress. Some potential ways forward are suggested. Thirdly, this article identifies the close relation between a model’s epistemic complexity and its interpretability, as introduced in the context of responsible AI. This clarifies in which sense—and to what extent—the lack of understanding of a model (black-box problem) impacts its interpretability in a way that is independent of individual skills. It also clarifies how interpretability is a precondition for a plausible assessment of the reliability of any model, which cannot be based on statistical analysis alone. This article focuses on the comparison between traditional scientific models and DNN models. However, Random Forest (RF) and Logistic Regression (LR) models are also briefly considered.
@@ 1 Introduction
First, we cannot completely exclude the possibility that successful DNN algorithms are over-fitting the collections of datasets that are used to train and test them (Recht et al., 2018). In fact, because of the difficulty of collecting good quality labelled data, very few very popular datasets (in particular CIFAR-10, Krizhevsky et al. (2014), ImageNet, Russakovsky et al. (2015) and a few others) represent the unique benchmark for the majority of the research work on DNN. This challenges one of the key statistical assumptions of all ML methods, namely that the parameters are set independently of the test data. In principle, nothing about the test data should be known in the design and training phase. In practice, train-test contamination can occur in subtle ways, even if we don’t directly use test data for training (Kapoor and Narayanan, 2023). Strictly speaking, if some data have been used in an article that we read before designing our new DNN architecture, those data should not be used to test our new architecture, because they might already have influenced our design choices indirectly. In practice, it can be hard to adhere to this requirement rigorously.
Secondly, successful applications of ML methods are much more likely to be published than unsuccessful ones. While publication biases (Dickersin et al., 1987) affect all types of scientific research, one might expect a stronger impact on ML research, because ML models are less likely to bring valuable insight when they fail. In fact, we have less compelling reasons to believe that they must succeed, which is related to the lack of full understanding of the mechanism that allows a DNN model to learn some features and ignore others.
Furthermore, assessment of confidence levels of ML predictions is notoriously difficult, especially in the case of DNN. The best evidence of these difficulties is provided by so-called adversarial examples (Szegedy et al., 2014): images that are misclassified with very high confidence by the DNN classifier, although they are humanly indistinguishable from other images that are correctly classified. Adversarial examples are much more pervasive in DNN than initially expected (Carlini and Wagner, 2017). The problem is not the occurrence of misclassification itself, but the very high confidence (easily $>99\%$) quoted by the classifier. This is clearly not a reliable error estimate [^1] It also does not help to say that adversarial examples have negligible probability among natural images, because such statement cannot be made quantitative, as we lack a suitable characterisation of the space of possible natural images. Moreover, adversarials have been also reproduced from natural images (Hendrycks et al., 2019)..
@@
ML methods in general, and DNN in particular, are being used successfully in many contexts where assessing their expected errors is not necessary. This applies, for example, to all those cases where the ML method improves the search efficiency of candidate solutions that can be subsequently verified by other means (Duede, 2023). In these cases, ML methods are used in the context of discovery, and not in the context of justification. Examples cover a vast range of applications, spanning material discovery, drug discovery, predictive maintenance, fraud detection (Baesens et al., 2015), code suggestion (Brown et al., 2020), protein folding (Jumper et al., 2021) and many others. In those cases, assessing the reliability is the responsibility of the independent check, while the ML method merely improves the efficiency of the overall process. However, there are also applications where independent checks are not practical (e.g. for safety-critical real-time systems (Buttazzo, 2022)) or cases where it is crucial to estimate the amount of missed solutions (e.g. compliance applications). In those cases, assessing the reliability of the ML method is very important.
This paper addresses the reliability of DNN methods from a fundamental epistemological point of view. It is very important to combine this perspective with a purely statistical one because even traditional science does not offer an absolute (assumption-free) guarantee that its predictions have specific probabilities. So, we must understand to what extent DNN models rest on similar grounds as those employed by more traditional scientific disciplines and to what extent they might suffer from fundamentally different problems.
@@
The topic of this article falls at the intersection of the theory of complexity (Franklin and Porter, 2020; Zenil, 2020a; Li and Vitányi, 2019; Hutter, 2007), DNNs (Goodfellow et al., 2016; LeCun et al., 2015), responsible AI (MacIntyre et al., 2021; Eitel-Porter, 2021) and the epistemology of ML, which has attracted considerable attention recently (see e.g. Desai et al. (2022); Leonelli (2020)). As opposed to many works on this topic (but similar to Kasirzadeh et al. (2023); Zenil and Bringsjord (2020); Grote et al. (2024) among others), this article stresses the important peculiarities of DNNs within ML methods. The present approach has some similarities with the one adopted in (Watson and Floridi, 2021). Some of the most important differences are: (a) the present focus is on reliability, that naturally leads to consider global, rather than local interpretability; (b) for the same reason, the relevance defined in (Watson and Floridi, 2021) is less applicable and it is not considered here; (c) finally, Watson and Floridi (2021) consider a subjective notion of simplicity, while we focus on the hardcore complexity that no human can reduce, regardless of language and individual skills. This article investigates the objective foundations of reliability assessment (i.e. based on well-defined assumptions). Hence, it is largely unrelated to the literature that investigates the sociological basis for the trust in ML models, or analogies between ML and human behaviours (see e.g., Duede (2022); Clark and Khosrowi (2022); Tamir and Shech (2023)). Further comparisons with the literature are provided in the main text.
@@ 2 Assessing the reliability of model predictions in science and ML
To assess the reliability of any model we must be able to estimate the uncertainty of its predictions, in some precise and useful probabilistic sense. To start, it is convenient to distinguish between statistical and systematic uncertainties. Traditionally, statistical uncertainties are defined as those error sources which have a known statistical distribution (Bohm and Zech, 2017) [^2] The ML literature adopts sometimes similar but slightly different classification between statistical (aka random, aleatoric or ‘data’) uncertainties on one side and systematic (aka epistemic or model) uncertainties on the other (Hüllermeier and Waegeman, 2021; Gawlikowski et al., 2021). Here we stick to the traditional one, because it clarifies the role of the model assumptions, which is crucial in our discussion.. These uncertainties can be safely analysed via statistical methods. Systematic uncertainties are all the others: they may stem from systematic distortions in the measurement devices, from imbalances in the data selection, from inaccurate models, from approximation errors, or from inaccurate parameter fitting. As emphasised in Bohm and Zech (2017), even random noise with unknown statistical properties must be classified as a systematic effect [^3] Note that probability distributions without finite mean and variance, for which the central limit theorem does not apply, cannot be characterised empirically in most cases. Given the ubiquity of, e.g., self-similar phenomena (Barenblatt, 1996) it is clear that one cannot certainly assume that every noise generates statistical uncertainties.. Statistical uncertainties can be very difficult to estimate in practice, but the process is conceptually clear. They are known-unknown, as they are expected by the model. Systematic uncertainties are unknown-unknown and they can be estimated only by enriching the model assumptions. This raises the question of which assumptions are acceptable to assess our model assumptions?
@@ 2.1 Sources of errors
It is useful to distinguish four different sources of errors and see how they affect ML and TS models in different ways. Errors may stem (a) from data measurements, including the labels of training data; (b) from the model, which may not faithfully describe the actual phenomena in scope; (c) from any approximation that may be applied to the model to derive further predictions; and (d) from fitting the parameters that are left free by the model [^5] Type (a) and (d) may have both statistical and systematic contributions, while type (b) and (c) are purely systematic. [^6] Models with free parameters are, strictly speaking, not yet specific enough to enable the derivation of predictions. They can be considered model-classes, but we simply use the term “model” so as not to burden the notation..
@@
Sources (b) and (d) are the most interesting ones when comparing TS and ML models. Model errors (b) can be reduced by extending the basic (ideal) model with further elements designed to model the deviations from the basic model. However, this enrichment comes at the expense of more complex assumptions and more parameters to be fitted (generating more uncertainties of type (d)).
Source (d) refers to errors in the determination of the optimal parameters, even when the model might provide a correct representation of the phenomena. This may happen because of limited data, noisy data (source (a)), an inaccurate minimisation algorithm or an inexact parameter fitting process. Error source (d) has a statistical component, which is propagated from noisy data (a) and can be analysed statistically (along the lines of e.g. Marshall and Blencoe (2005) and references therein). However, source (d) also has systematic components, due to the complexity of determining the optimal parameters and also due to the unknown statistical properties of the data777for some classes of noise sources, collecting sufficient data can be completely hopeless in practice (see, e.g., de Forcrand (2009)) or even in principle (see e.g., Gabaix (2009)).. Moreover, increasing the number of parameters requires, in general, an exponential increase in the amount of data required for their determination [^8] An exponential increase is expected, in the generic case, according to the curse of dimensionality (Giraud, 2021).. In conclusion, also estimating (d) requires making suitable assumptions.
@@ 2.2 The illusion of predictions without assumptions
The simple reason why this idea is fallacious is the well-known (but often overlooked) underdetermination of the theory by the data (Duhem, 1954; Quine, 1975; Stanford, 2021). No matter how much data we have collected, there are always infinite different models that fit those data within any given approximation. Moreover, for every conceivable prediction, there exist infinite models that still fit all the previous data and also deliver the desired prediction. Data alone cannot enforce any conclusion.
@@
Another frequent fallacy is the idea that we should only accept assumptions that have been tested. But how much and under which conditions should they be tested? Newton’s laws were tested extensively under all conceivable circumstances for over two centuries before discovering that they do not always hold. There is no way to fully test our assumptions (or to define their scope of applicability) in a future-proof way. They may fail in contexts that we cannot currently foresee.
@@ 3.1 Bayesian Error Estimates
However, such an argument is based on the unrealistic idea that the model is fixed, while we can rather easily collect more data. This is not what usually happens in real applications. In many cases, instead, collecting valuable data can be very hard, while producing new models is often much easier. This is the case in many TS domains, where planning a new experiment might take decades, while new models are published every day. It is the case also in most ML applications, where collecting labelled data can be extremely labour intensive, while new ML models are produced at the rate of millions per second by the ML algorithm. Under these conditions, there is no guarantee that the dependency on the priors will be sufficiently suppressed by the time the model is used.
@@
Increasing the number of parameters might look like a way to eliminate the need of changing the model. But, the Doob and Berstein - von Mises theorems do not apply when the number of parameters is of the same order as the size of the dataset or more (Johnstone, 2010; Sims, 2010). Extensions exist for high-dimensional problems (Banerjee et al., 2021) and even infinite-dimensional (aka nonparametric) ones (Ghosal and Van der Vaart, 2017). However, those extensions rely on the assumption that most parameters are very close to zero, which is achieved via LASSO priors or equivalent suppression mechanisms (Banerjee et al., 2021). Note that this is a much stronger assumption than requiring that all the relevant parameters live in some low-dimensional subspace of a high-dimensional model. A parameter suppression mechanism imposes a sparse realisation of the given parametrisation. As a result, even in the asymptotic limit of infinite statistics, the posteriors depend on the choice of the parametrisation, similar to the case of low-dimensional models. In other words, nonparametric models are not a way to eliminate model dependency.
The limitations of nonparametric models confirm the conclusions of Sec. 2.2 about the impossibility of estimating errors that take into account all possible models. The Bayesian approach is an excellent framework to discuss the probability of a model within a given class (say $C$) of models, but it cannot justify the selection of one class $C$ of models over other options.
In principle, the class $C$ should include a model that describes well all past and future data (what is sometimes called a true model [^10] Note that a true model defined in this way might or might not exist or might not be unique in the given parametrization.) [^11] To be precise, if the goal is to assess the reliability of a model prediction, then the class $C$ should contain a true model, as stated in the main text. If the goal is to test a model, then the requirement applies to the background models to be used in combination with the model to be tested (i.e. the background part should include a true model). See Sec. 4.1 for more details on background models.. However, we cannot know which are the true models, of course, and we have to resort to other criteria.
The ambiguity in the selection of the class $C$ applies to both TS and DNN models. However, in practice, the scientists tend to agree on which other models should be included to assess the errors of a given TS model. Ideally, they should include all the models that might provide legitimate alternative descriptions of the relevant phenomena, because they offer different compromises between accuracy (for different phenomena) and non-empirical epistemic virtues. These are sometimes referred to as state-of-the-art models, but the class is not well defined unless we identify the appropriate non-empirical epistemic virtues. We will come back to this point in Sec. 4.2, where we consider a definition of state-of-the-art based on the simplicity of the assumptions (see also App. B). We will also see that identifying a similar class for DNN models is significantly more challenging.
@@ 3.4 Predictions are NOT all you need
Can’t we just use the success rate of the test dataset as an estimate of the expected prediction error? Intuitively, one expects that novel predictions on unseen data do provide a meaningful test. But, from a statistical point of view, the success rate corresponds to the simplest frequentist statistical estimator, with the limitations already discussed in Sec. 3.2. From a philosophical point of view, the idea that successful tests confirm a model is fraught with riddles (Crupi, 2021; Huber, 2014). In fact, for example, we cannot be sure that the future data will conform to the same distribution of the past data. One might argue that this corresponds to a change of domain of application and the model cannot be blamed for that. But we cannot define the scope of applicability of any model in a way that is future-proof: the model may fail in circumstances that we currently cannot even imagine. All the attempts to quantify the degree to which the evidence confirms a model have shown the need to rely on extra assumptions which are themselves hard to justify (Huber, 2014). For these reasons, predictions alone—no matter how impressive—are never sufficient to warrant trust in a model, and Likelihood—by itself—is never sufficient to determine model selection. Other non-empirical aspects are also important. But which ones?
These problems also apply to TS models, but are more evident for DNN models whose persuasive power relies relatively more on their empirical success than on their theoretical virtues. Although DNN successes are impressive, it is impossible to derive any quantitative level of confirmation from them without additional assumptions. For example, published results certainly suffer a positive selection. Kaggle competitions (Goldbloom, 2015) are very interesting also from a philosophical point of view because they provide a controlled framework for assessment. However, they also cannot be used directly to confirm either general DNN architectures or specific DNN models, because submissions are strongly biased in terms of the approach used and the competencies of the participants. The fact that some successful DNN architectures keep being successful over time is certainly convincing evidence that they can learn something valuable. However, also in those cases, we cannot exclude that the DNNs have actually learned unwanted features that happened to have a strong spurious correlation with the labels (see, e.g., Xiao et al. (2021) for an example of this phenomenon). This suggests that DNN predictions are only valuable when they stem from a valuable model. What characterises a valuable model, both in TS and in ML? This question is addressed in more depth in Sec. 4.